OpenStudy (anonymous):

i really don't understand these..2x^2+5x-3 i know some of you helped but i don't understand

Parth (parthkohli):

Again, this is a quadratic expression.

Parth (parthkohli):

Do you NOW wanna know how to do these?

OpenStudy (anonymous):

yes...i have to do this by tonight so plz help

Parth (parthkohli):

All quadratic expressions after grouping are expressed as \(\Large \color{purple}{\rightarrow (zx + a)(x + b) }\)

Parth (parthkohli):

Now, let's see it in detail.

Parth (parthkohli):

Also, quadratic expressions are in the form --> ax^2 + bx + c

OpenStudy (anonymous):

What do you want to do with these?

OpenStudy (anonymous):

can you explain step by step how to do them? i have alot more like these..

Parth (parthkohli):

To factor, ac must be equal to the numerical coefficient of the middle term.

Parth (parthkohli):

http://www.youtube.com/watch?v=eF6zYNzlZKQ This guy is magical.

jimthompson5910 (jim_thompson5910):

2x^2+5x-3 2x^2-x+6x-3 (2x^2-x)+(6x-3) x(2x-1)+(6x-3) x(2x-1)+3(2x-1) (x+3)(2x-1) So 2x^2+5x-3 completely factors to (x+3)(2x-1)

jimthompson5910 (jim_thompson5910):

Notice that 2*(-3) = -6. Two numbers that multiply to -6 and add to 5 are -1 and 6. So we break up 5x into -x+6x. From there, you factor by grouping.

OpenStudy (anonymous):

thanks

Directrix (directrix):

A little more on factoring quadratics: I learned the "bust the b" technique. 4 x² -1x + 18 a = 4 and c = 18 and b = -1. The task on "bust the b" is to find numbers that mutiply to a*c AND add to b. In this case, numbers that multiply to 4*18 and add to -1. 14 * 18 breaks aparts into the product of prime factors : 2*2*2*3*3 Look at those five factor of 14 * 18 and try to put them in the form of 2 numbers that differ by -1. 2*2*2 and 3*3 appear to differ by 1. That would be 8 and 9. "b" has been "busted into 8 and - 9 whose sum is -1 which equals "b" in this problem. 4x - x+ 18 = 4 x² + 8 x - 9 x - 1. The "b" has been busted. Factoring by grouping comes next. 4 x² + 8 x - 9 x - 18 = 4x ( x + 2) - 9 ( x + 2 ) = --> On this step, (x + 2) is the common factor of the expression. [ (x + 2) ] ( 4x - 9 ) = ( x + 2 ) ( 4x - 9)